OpenStudy (anonymous):

prove that, for all x>1, 2*sqrt(x) >3-1/x @Mathematics

OpenStudy (amistre64):

could you solve it using and a > b case, and a=b case and an a<b case?

OpenStudy (amistre64):

spose a<b, but that a,b > 1 2sqrt(a) > 3 -1/b 4a > 9 -6/b -1/b^2 b^2 4a -9b^2+6b+1 > 0 b^2(4a -9) + 6b + 1 > 0 b = -6+- sqrt(36-4(1)(4a-9)) /2 b = -6+sqrt(36-16a +36)) /2 b = -6+2sqrt(18-4a) /2 b = -3+sqrt(18-4a) i hate proofs, really i do .....

OpenStudy (amistre64):

any particular strategy?

OpenStudy (anonymous):

it wants us to use the first and second derivative to prove it using limits at infinity

OpenStudy (asnaseer):

let \[a = 2\sqrt {x}\]\[b = 3 - 1/x\]then let \[y = a ^{2} = 4x\]\[z = b ^{2} = (3 - 1/x)(3 - 1/x) = (3x - 1)(3x - 1)/x ^{2} = (9x ^{2} - 6x + 1)/x ^{2}\]\[ = 9 - 6/x + 1/x ^{2}\] when x = 1, y = 4 and z = 9 - 6 + 1 = 4 as x increases, z tends to 9 as both 1/x and 1/x^2 tend to zero so, for x > 1, y > z

OpenStudy (anonymous):

thanks for your help!

OpenStudy (amistre64):

well, the first derivative would tell us how fast it changes and teh second tells us about the speed at which the speed is changeing

OpenStudy (amistre64):

f(x) = 2sqrt(x); g(x)=3 -1/x prove that f(x) > g(x), for all x>1

OpenStudy (amistre64):

f'(x) = 1/sqrt(x) ; g'(x) = 1/x^2 f''(x) = -1/2sqrt(x^3) ; g''(x) = -2/x^3 if i did those right

OpenStudy (amistre64):

infinity in the bottoms would both equate to 0 for both derivatives so i got no idea how that helps

OpenStudy (amistre64):

lim(x->inf) f(x) = inf lim(x->inf) g(x) = 3 that might be useful

OpenStudy (anonymous):

thank you @amistre64!

OpenStudy (amistre64):

youre welcome, good luck ;)